Operation Summary for November 11, 2011 - November 14, 2011 |
NuMI reported receiving 7.53E18 protons on target for the period from 11/7/11 to 11/14/11. |
|
Friday November 11 The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, with Neutron Therapy Facility (NTF) personnel preparing to treat two patients later in the morning, with the Muon Ring power supplies (formerly the Antiproton Source) turn on in progress, and with the NuMI and MiniBooNE experiments taking beam. During the day shift, Operations established beam to the Meson FTBF experiment T-978 (CALICE) around 8 AM. Operators reported at 8:12 AM that they had turned on the Muon rings, Pretarget, and Prevault power supplies. A Muon person began a one-shot study at 9:46 AM, followed by a Debuncher Transverse Schottky tune measurement, and then a resumption of the one-shot study. Booster personnel began a one-hour study at 9:47 AM. At 10:01 AM, the NuMI beam permit tripped off due to bad beam positions at the target; everything reset. Operators turned off the MiniBooNE beam at 10:28 AM for about a 30-minute access to the MI-12 service building. A Linac person completed changing the I- Source’s hydrogen bottle at 11:21 AM. An operator completed a tune of the 750 KeV line at 12:27 PM. At 1:23 PM, MI personnel conducted a Recycler chromaticity measurement. An operator completed a tune of Booster injections at 3:12 PM. The Muon Ring studies ended and operators prepared the area for access at 3:56 PM. During the evening shift, the MI Recycler studies ended at 5:24 PM. Operators disabled beam to T-978 at 7:07 PM. Operators prepared the Meson M01 enclosure for an access to check the level of water in a sump pump; it was okay. At 8:11 PM, operators reported a MI-60 vacuum burst that was coincidental with a NuMI kicker (KPS6N) trip; operators contacted and called in MI and Vacuum personnel. Operators prepared MI and MI-8 for an access at 9:42 PM. Saturday November 12 The midnight shift began with MI vacuum leak investigations. During the midnight shift, at 1:42 AM, a Vacuum person reported at finding as big leak on the flange between the 604 and 605 location; Vacuum technicians replaced a flange gasket. At 5:18 AM, Vacuum technicians reported that the vacuum was recovering slowly in the 604 region, but they found another leak between 603 and 604. Vacuum techs reported at 7:24 AM that they had replaced another flange gasket in the 603 region and had started the vacuum pump down. During the day shift, a Mechanical Engineer reported at 11:13 AM that the MI vacuum was stable. Operators opened the beam valves and began the MI checkout. Operations established beam to MI at 11:54 AM. Operators reported a P2 transfer line power supply (I:HVF12) was holding off the beam permit; operators called in Power Supply personnel to investigate. Operations established beam to NuMI and MiniBooNE by 1:08 PM. Operations established beam to T-978 at 3:13 PM. At 3:36 PM, Power Supply personnel reported that they had found a LCW leak on a SCR module; they ended up replacing two SCR modules. During the evening shift, Operations disabled beam to T-978 at 7:10 PM. The NuMI beam permit tripped off at 7:29 PM; it reset. Operators reported at 10:31 PM that a Linac Low Level RF node for an RF station (LRF2) and an ion pump (LIP742) was not responding; operators reported this to a Linac person who said they would need two hours of down time to complete a reboot. Sunday November 13 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, at 4:06 AM, operators reported that the MI-60 BPM node was not responding; operators rebooted the system. At 6:28 AM, operators reported that the MI/NuMI Lambertson (LAM61) was causing a beam permit to trip off; operators called in power supply personnel. Operations put NuMI in standby and increased the beam intensity to MiniBooNE. During the day shift, Operations reestablished beam to NuMI at 9:02 AM. Operations established beam to T-978 at 9:04 AM. Power supply personnel reported at 9:09 AM that they had replaced a SCR and a firing card for LAM61. At 2:13 PM, a NuMI BPM (NBPM65) stopped responding; operators rebooted the system. During the evening shift, at 4:17 PM, the NuMI beam permit tripped off when the current for a quadrupole magnet's power supply (QF114) drifted low: everything reset. Operations disabled beam to T-978 at 7:11 PM. Operators rebooted a NuMI BPM VME at 10:30 PM. Monday November 14 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, Operators rebooted a NuMI BPM VME at 12:57 AM. At 5:06 AM, the NuMI beam permit tripped off due to RF gap faults; it reset. At 5:33 AM, the NuMI beam permit tripped off due to a BPM not polling; a reboot fixed the problem. Operations turned off beam to MiniBooNE at 5:56 AM to replace a water skid pump later in the morning. Operators completed a 750 KeV line tune at 6:20 AM. At 6:45 AM, BRF17 and BRF18 tripped off on a high vacuum pressure indication; everything reset. Operations established beam to T-978 at 7:02 AM. Operators rebooted the MI-30 BPMs at 7:55 AM because they had stopped responding; everything reset. Operations Coordinator (OpsCo) The OpsCo approved of the plans by Booster and MI personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 11/14/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |